# Using neuroevolution for designing soft medical devices

### v1

@@ 5 Conclusions
This research studied the suitability of NEAT and HyperNEAT as a generator of catheter actuator morphologies. Their capabilities were compared against an implementation of AFPO, a popular multi-objective optimisation approach. The performance analysis was done utilising three metrics: (i) analysing the general performance to find actuators with the maximum displacement possible in the $yz$ plane; (ii) testing the robustness of those actuators exhibiting the maximum displacement reached (i.e., the fittest actuators); and (iii) determining the fittest actuators with the minimum number of voxels possible. For this analysis, 20 different actuators were generated (i.e., 20 experimental runs were conducted) by each approach.
@@
Results suggest that neuroevolution-based approaches are generally more capable of producing robust catheter actuator morphologies than AFPO due to the genetic operators (i.e., crossover and mutation) induced a broader exploration in the search space, leading to the discovery of fitter individuals than those discovered by AFPO. In addition, the actuator produced by NEAT could outperform the one generated by HyperNEAT. Arguably, the design of the substrate utilised during experimentation affected the performance of HyperNEAT.
@@
Furthermore, the number of voxels plays a crucial role in bending tasks; the fittest actuator found by AFPO exhibited significantly fewer contractile voxels; however, its performance was poor. On the other hand, the fittest actuator generated by NEAT and HyperNEAT showed at least twice the number of voxels of the actuator generated by AFPO. Their performance, nevertheless, overcomes the one from the actuator generated by AFPO. Arguably, the striped no-voxel pattern exhibited by NEAT and HyperNEAT actuators helped to enhance the bending movement.
@@
Considering the results and insights gathered from this research, future work may include the addition of more periodic activation functions, such as cosine and tangent. Following that tactic will help to induce more diverse morphological patterns produced by CPPNs. Another avenue of future work consists of a broader exploration of the number of hidden layers and neurons that may lead to better performance under HyperNEAT. In order to improve the design of the substrate, an approach called ES-HyperNEAT, focused on evolving the location of every neuron and the pattern of weights among them, can be implemented [47]. Another crucial aspect of future work will be focused on evolving controllers for the fittest catheters actuators found in this research.

### v2

@@ 5 Conclusions
This research studied the suitability of NEAT and HyperNEAT as generators of soft actuator morphologies. Their capabilities were compared against an implementation of AFPO, a popular multi-objective optimisation approach. The performance analysis was done utilising three metrics: (i) analyzing the general performance to find actuators with the maximum displacement possible in the $yz$ plane; (ii) testing the robustness of those actuators exhibiting the maximum displacement reached (i.e., the fittest actuators); and (iii) determining the fittest actuators with the minimum number of voxels possible. For this analysis, 20 different actuators were generated (i.e., 20 experimental runs were conducted) by each approach.
@@
Results suggest that neuroevolution-based approaches are generally more capable of producing robust catheter actuator morphologies than AFPO due to the genetic operators (i.e., crossover and mutation) induced a broader exploration in the search space. In addition, the actuator produced by NEAT could outperform the one generated by HyperNEAT. Arguably, the design of the substrate utilised during experimentation affected the performance of HyperNEAT.
@@
Furthermore, the number of voxels plays a crucial role in bending tasks; the fittest actuator found by AFPO exhibited significantly fewer contractile voxels; however, its performance was not superior when considering the multitude of possible controller strategies. On the other hand, the fittest actuator generated by NEAT and HyperNEAT showed at least twice the number of voxels of the actuator generated by AFPO. Their performance, nevertheless, overcomes the one from the actuator generated by AFPO. Arguably, the striped no-voxel pattern exhibited by NEAT and HyperNEAT actuators helped to enhance the bending movement. Although, the designs produced by the proposed NE-based methodologies were not validated in vitro, this should not hinder their applicability or lessen their better performance demonstrated when compared with AFPO. The validation of AFPO applied on Voxelyze has been previously proved after fine-tuning the environment simulation parameters [34]. The selection of materials to build the actuator are considered as the material parameters inputs to Voxelyze, while the manufacturing process of artificial material should not hinder the generalization of this methodology.
The fact that the fitness function of the morphologies considered in the evolutionary optimisation methodology is provided by Voxelyze, a voxel-based physics simulator, may give rise to some concern around the translation of designs to real-world applications through concurrent manufacturing processes. Voxelyze utilises cube elements as building blocks and, thus, the morphologies produced may be predominately comprised by orthogonal boundaries; a characteristic that is not ideal for real-world applications. Namely, soft material (i.e., plastic melted through a 3D printer nozzle) or biohybrid material (i.e., living cells that are subject to tissue adherence/contraction) can not formulate perfectly right angles in the micro-scale. Nonetheless, the implementation of the evolutionary optimisation through indirect representations (i.e., CPPNs) can alleviate part of this limitation. These representations are scale free. Consequently, the anatomical resolution of the morphology may be increased without any further tests of the computational demanding evolutionary optimisation. The enhanced resolution may develop convex morphologies that would appear as orthogonal in lower resolutions [34]. The alleviation of right angles can be further enhanced through the use of appropriate activation functions on NEAT algorithm. On the other hand, this is not true for all cases. Therefore, an implementation of filtering the morphologies for manufacturing compatibility can follow the morphology generator proposed here, in the automated design pipeline envisioned. These filters will remove morphologies that are not possible or trivial to build with the available manufacturing techniques. However, the implementation of these filters are considered as an aspect of future work.
@@
Considering the results and insights gathered from this research, future work may include the addition of more periodic activation functions, such as cosine and tangent. Following that tactic will help to induce more diverse morphological patterns produced by CPPNs. Another avenue of future work consists of a broader exploration of the number of hidden layers and neurons that may lead to better performance under HyperNEAT. In order to improve the design of the substrate, an approach called ES-HyperNEAT, focused on evolving the location of every neuron and the pattern of weights among them, can be implemented [58].
Moreover, a more robust performance comparison against other NE-based approaches will be conducted. For instance, Evolutionary eXploration of Augmenting Memory Models (EXAMM), which is capable of evolving ANNs (including recurrent ANNs) through a wide variety of memory structures [53, 72]. Taking into account that real-world optimisation problems can be time-constrained, the required computational time of further methodologies studied will be a key factor. Furthermore, the utilisation of surrogate-model assisted optimisation [57] will be considered, to minimize the need for computationally demanding physics simulators. Finally, another crucial aspect of future work will be focused on evolving controllers for the fittest catheters actuators found in this research.